# Dynamics of “Spontaneous” Topic Changes in Next Token Prediction with Self-Attention

### v1

@@ 5 Explaining topic shifts
The formation of new SCCs when combining datasets suggests that the highest priority SCC for some input sequences may increase in size in this new setting. This also suggests that topic shifts may arise from ambiguity within an input sequence rather than a straightforward change in topic. In our analogy on the oracle, gaining knowledge of both Topic A and Topic B might cause a conversation to be naturally followed within Topic A or also outside Topic A. Thus, we introduce the following definition to characterize this phenomenon:
@@
Given DSETa and DSETb generated from two different topics $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, and a sequence $\bf X$ that belongs to $\mathbb{T}_{a}$. The weight matrix $\tilde{\bf W}_{ab}$ changes topic $\mathbb{T}_{a}$ for sequence ${\bf X}$ if $\tilde{\bf W}_{ab}$ does not keep topic $\mathbb{T}_{a}$ for ${\bf X}$ and $\bf X$ is not ambiguous in $\mathbb{T}_{ab}$ with respect to $\mathbb{T}_{a}$.
Following the above definition, ${\bf W}_{ab}$ for the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$, from Figure 2 (bottom), can be considered as change of topic.
@@
Consider a sequence ${\bf X}$ that belongs to $\mathbb{T}$. We define $\dot{\mathcal{G}}^{(\bar{x})}({\bf X})\in\mathcal{G}^{(\bar{x})}$ as the highest priority SCC for ${\bf X}$ in $\mathcal{G}^{(\bar{x})}$ such that $\forall{\bf x}_{i}\in\dot{\mathcal{G}}^{(\bar{x})}({\bf X})$ and ${\bf x}_{j}\in\mathcal{G}^{(\bar{x})}$ we have $(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}^{(\bar{x})}$.
@@
There are two implications of this theorem. First, as the input sequence length increases sufficiently, the likelihood of topic changes vanishes. Second, increasing $l$ leads to possibly higher overlap between topics, meaning that the probability of satisfying the necessary conditions for a topic change decreases, effectively creating a bound on the frequency of topic changes. In practice, consider the oracle analogy: if the oracle is following a sufficiently long conversation on a specific topic, it becomes exceedingly unlikely to shift topics. Similarly, as topics A and B become more interconnected, this increased ambiguity does not lead to more topic changes; rather, it may reduce their occurrence. This contrasts with human cognition, where longer conversations and greater inter-connectivity of knowledge increase the likelihood of spontaneous shifts.
(a) Effect of increased sequences.
(b) Effect of increased overlap.
Figure 4: The proportion of keep topic, ambiguity, and change topic as (a) the test sequence length increases and (b) the number of edges in generated TPGs increase.
To illustrate Theorem 5, we conduct simulations following the same implementation settings as described in Section 4. We approximate $\tilde{\mathbf{W}}_{a}$ and $\tilde{\mathbf{W}}_{ab}$ as the results obtained after $\tau=8000$ iterations of Algo-GD. We quantify the proportion of test sequences in which $\tilde{{\bf W}}_{ab}$ keeps $\mathbb{T}_{a}$ (keep topic), proportion of ambiguous sequences in $\mathbb{T}_{ab}$ (ambiguity) and proportion of times in which $\tilde{{\bf W}}_{ab}$ changes topic (change topic). First, we explore the effect of longer sequences by varying the length $T$ of the test sequences $\bf Z$. We increase $T$ from 4 to 512. Figure 4 illustrates how the proportion of change topic decreases as $T$ increases. Second, we investigate the effect of overlapping topics by generating TPGs with an increased number of edges $L$. Intuitively, a higher $L$ results in greater overlap between TPGs from different topics. We vary $L$ from 4 to 18. Figure 4 demonstrates that as $L$ increases, ambiguity increases, while proportion of change topic decreases. These two findings contrast with expectations derived from human cognition but align with the result of Theorem 5. Lastly, among the 85000 test sequences generated for these experiments, 99.98% satisfy Theorem 4 (i.e., topic changes occur when a low-priority token appears more frequently than high-priority tokens). The remaining 0.02% mismatched cases were solely due to minor approximation discrepancies in the attention softmax. These results validate Theorem 4 (see specific experimental details in the Appendix).
@@
Training and generalization of Transformer. (1) Properties of Self-attention and Softmax. The self-attention mechanism employs the softmax function to selectively emphasize different parts of the input. Gu et al., (2024), Goodfellow et al., (2016), and Deng et al., (2023) underscore the pivotal role of the softmax function in shaping attention distributions, influencing how models process and prioritize information within input sequences. Bombari and Mondelli, (2024) examined the word sensitivity of attention layers, revealing that softmax-based attention-layers are adept at capturing the significance of individual words. Despite these strengths, recent researches have highlighted limitations of the softmax function (Saratchandran et al.,, 2024; Deng et al.,, 2023). (2) Optimization in attention-based models. Additionally, recent researches interpret Transformer models as kernel machines, akin to support vector machines (SVMs), with self-attention layers performing maximum margin separation in the token space (Tarzanagh et al.,, 2023; Ataee Tarzanagh et al.,, 2023; Li et al., 2024b, ; Julistiono et al.,, 2024). (3) Chain-of-Thought and In-Context Learning. Moreover, transformers exhibit remarkable abilities in generalization through in-context learning (ICL), where models effectively learn from contextual cues during inference (Brown et al.,, 2020; Xie et al.,, 2022; Olsson et al.,, 2022). Chain-of-Thought (CoT) prompting (Wei et al.,, 2022; Zhou et al.,, 2022; Shao et al.,, 2023; Li et al., 2024a, ) enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers. (4) Improvement efficiency of transformers. Recent advancements aim to improve the computational efficiency of transformers (Kitaev et al.,, 2020; Choromanski et al.,, 2021; Sukhbaatar et al.,, 2019; Wang et al.,, 2020), ensuring their viability for large-scale deployment while maintaining or enhancing their representational capabilities.
@@ 6 Discussion
Our findings on a simple single-layer attention model provide valuable clues about the fundamental differences in conversational dynamics between humans and machines. Unlike humans, who naturally shift topics through curiosity or contextual association, the model adheres to predictive patterns, mirroring LLMs’ tendency to stay focused on a single thread. This suggests that spontaneous topic shifts, a hallmark of human conversation, may require mechanisms beyond predictive pattern recognition. These insights raise philosophical questions about the nature of human-like interaction and highlight the need for adaptability and creativity in advancing conversational AI.
Future research will focus on how real-world LLMs perform changing topics, where the interplay of diverse and evolving contexts presents unique challenges. It would also be interesting to investigate the connection between LLM and human cognition in spontaneously changing topics. Additionally, the recurring patterns of topic retention identified in this work open new pathways for applications. Harnessing these behavioral signatures could enhance changing topics detection in conversational AI, enrich content generation by anticipating and adapting to topic transitions, and refine model interpretability. Pursuing these directions could significantly advance both the theoretical landscape and practical utility of LLMs.
@@ A.4 Proof of Theorem 5
Appendix B Detailed simulation studies
B.1 Simulation process
@@
Next token prediction. To differentiate the input sequence length of the testing data from that of the training data, we introduce $T_{test}$. TPGs based on the training dataset DSETa are utilized to generate test datasets consisting of 100 sequences from $\mathbb{T}_{a}$ per epoch. Specifically, the last token ${\bf x}_{T_{test}}$ of the test input sequence is randomly selected from $K$ tokens (i.e. ${\bf x}_{T_{test}}={\bf e}_{k}$) and the remaining input tokens are randomly chosen based on the SCCs of $\mathcal{G}^{k}_{a}$, where tokens with higher priority are assigned greater weights. For instance, in $\mathcal{G}^{4}_{a}$, tokens ${\bf e}_{1},{\bf e}_{2},{\bf e}_{4}$ are captured with the priority order ${\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$. The weights assigned to input tokens ${\bf e}_{1},{\bf e}_{2}$, and ${\bf e}_{4}$ are $0.4,0.4$, and $0.2$, respectively. It reflects that ${\bf e}_{1}$ and ${\bf e}_{2}$ are in the same higher-priority SCC, thus having greater weights compared to ${\bf e}_{4}$. Intuitively, tokens within the same SCC are more likely to co-occur than those from different SCCs. This approach enables the generated test input sequences to mimic real word relationships and reflect their contextual groupings. Following the generation of the test dataset from $\mathbb{T}_{a}$, the next tokens $\hat{y}_{\bf w_{a}}$ and $\hat{y}_{\bf w_{ab}}$ are predicted by Equation 1, with ${\bf W}_{a}$ and ${\bf W}_{ab}$ obtained from the last iteration. To reduce the potential numerical issues in the outputs, $\mathbb{S}({\bf X}{\bf W}\bar{{\bf x}})$ is rounded to three decimals, ensuring that tokens within the same SCC yield consistent softmax outputs.
B.2 Experiments results
@@ B.2.1 Simulation in Section 4
B.2.2 Simulation in Section 5
For Figure 4 (left), we predict next tokens for 5000 test sequences from $\mathbb{T}_{a}$ with $T_{test}=\{4,8,16,24,32,64,128,256,512\}$, while fixing $L=4$, $d=16$, $T_{train}=4$, and $K=10$. The proportion of each scenario with varying $T$ is illustrated in Table 1. For Figure 4 (right), we predict next tokens for 5000 test sequences (the sequence length is $T_{test}=20$) using models trained with $L=\{4,6,8,10,12,14,16,18\}$, $d=16$, $K=10$, and $T_{train}=4$. The proportion of each scenario with varying $L$ is illustrated in Table 2.
@@ B.2.2 Simulation in Section 5
B.2.3 Additional experiments
@@
B.3 Numerical analysis of each scenario in Figure 2

### v2

@@ 5 Explaining Topic Shifts
The formation of new SCCs when combining datasets suggests that the highest priority SCC for some input sequences may increase in size in this new setting. This also suggests that topic shifts may arise from ambiguity within an input sequence rather than a straightforward change in topic. In our analogy on the oracle, gaining knowledge of both Topic A and Topic B might cause a conversation to be naturally followed within Topic A or also outside Topic A. We introduce the following definition to characterize this phenomenon:
@@
Given DSETa and DSETb generated from two topics $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, and a sequence $\bf X$ that belongs to $\mathbb{T}_{a}$. The weight matrix $\tilde{\bf W}_{ab}$ changes topic $\mathbb{T}_{a}$ for sequence ${\bf X}$ if $\tilde{\bf W}_{ab}$ does not keep topic $\mathbb{T}_{a}$ for ${\bf X}$ and $\bf X$ is not ambiguous in $\mathbb{T}_{ab}$ with respect to $\mathbb{T}_{a}$.
In Figure 4, ${\bf W}_{ab}$ for the last input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$ can be considered as change of topic.
@@
Consider a sequence ${\bf X}$ that belongs to $\mathbb{T}$. We define $\dot{\mathcal{G}}^{(\bar{x})}({\bf X})\subseteq\mathcal{G}^{(\bar{x})}$ as the highest priority SCC for ${\bf X}$ in $\mathcal{G}^{(\bar{x})}$ such that $\forall{x}_{i}\in\dot{\mathcal{G}}^{(\bar{x})}({\bf X})$ and ${x}_{j}\in\mathcal{G}^{(\bar{x})}$ we have $(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}^{(\bar{x})}$.
@@
There are two implications of this theorem. First, as the input sequence length increases sufficiently, the likelihood of topic changes vanishes. Second, notice that increasing $l$ increases the probability of overlap between topics, and it also decreases the probability of satisfying the necessary conditions for a topic change, effectively creating a bound on the frequency of topic changes. In practice, consider the oracle analogy: if the oracle is following a sufficiently long conversation on a specific topic, it becomes exceedingly unlikely to shift topics. Similarly, as topics A and B become more interconnected, this increased ambiguity does not lead to more topic changes; rather, it may reduce their occurrence. This contrasts with human cognition, where longer conversations and greater inter-connectivity of knowledge increase the likelihood of spontaneous shifts.
(a) Length of sequences.
(b) Overlap between topics.
Figure 6: The proportion of keep topic (topic continuity), ambiguity (ambiguous sequence), and change topic (change of topic) as (a) the input sequence length increases and (b) the overlap between topics increases.
To illustrate Theorem 5, we conduct simulations following the same implementation settings as described in Section 4. We approximate $\tilde{\mathbf{W}}_{a}$ and $\tilde{\mathbf{W}}_{ab}$ as the results obtained after $\tau=8000$ iterations of Algo-GD. We quantify the proportion of test sequences in which $\tilde{{\bf W}}_{ab}$ keeps $\mathbb{T}_{a}$ (keep topic), proportion of ambiguous sequences in $\mathbb{T}_{ab}$ (ambiguity) and proportion of times in which $\tilde{{\bf W}}_{ab}$ changes topic (change topic). First, we explore the effect of longer sequences by varying the length $T$ of the test sequences $\bf Z$. We increase $T$ from 4 to 512. Figure 6(a) illustrates how the proportion of change topic decreases as $T$ increases. Second, we investigate the effect of overlapping topics by generating TPGs with an increased number of edges $L$. Intuitively, a higher $L$ results in an increase $l$ and a greater overlap between TPGs of different topics. We vary $L$ from 4 to 18. Figure 6(b) demonstrates that as $L$ increases, ambiguity increases, while proportion of change topic decreases. These two findings contrast with expectations derived from human cognition but align with the result of Theorem 5. Lastly, among the 85000 test sequences generated for these experiments, 99.98% satisfy Theorem 4 (i.e., topic changes occur when a low-priority token appears more frequently than high-priority tokens). The remaining 0.02% mismatched cases were solely due to minor approximation discrepancies in the attention softmax. These results validate Theorem 4 (simulation details in Appendix C).
@@
Training and generalization of Transformer. (1) Properties of Self-attention and Softmax. The self-attention mechanism employs the softmax function to selectively emphasize different parts of the input. Gu et al., (2024), Goodfellow et al., (2016), and Deng et al., (2023) underscore the pivotal role of the softmax function in shaping attention distributions, influencing how models process and prioritize information within input sequences. Bombari and Mondelli, (2024) examined the word sensitivity of attention layers, revealing that softmax-based attention-layers are adept at capturing the significance of individual words. Despite these strengths, recent researches have highlighted limitations of the softmax function (Saratchandran et al.,, 2024; Deng et al.,, 2023). (2) Optimization in attention-based models. Additionally, recent researches interpret Transformer models as kernel machines, akin to support vector machines (SVMs), with self-attention layers performing maximum margin separation in the token space (Tarzanagh et al., 2023a, ; Tarzanagh et al., 2023b, ; Li et al., 2024b, ; Julistiono et al.,, 2024). (3) Chain-of-Thought and In-Context Learning. Moreover, transformers exhibit remarkable abilities in generalization through in-context learning (ICL), where models effectively learn from contextual cues during inference (Brown et al.,, 2020; Xie et al.,, 2022; Olsson et al.,, 2022). Chain-of-Thought (CoT) prompting (Wei et al.,, 2022; Zhou et al.,, 2023; Shao et al.,, 2023; Li et al., 2024a, ) enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers. (4) Improvement efficiency of transformers. Recent advancements aim to improve the computational efficiency of transformers (Kitaev et al.,, 2020; Choromanski et al.,, 2021; Sukhbaatar et al.,, 2019; Wang et al.,, 2020), ensuring their viability for large-scale deployment while maintaining or enhancing their representational capabilities.
@@ 6 Discussion
Our exploration with a simple single-layer attention model reveals fundamental clues on the distinctions between human and AI conversational dynamics. While humans naturally shift topics through curiosity or contextual triggers, the model remains anchored to the context. These observations suggest that human-like spontaneity in topic changes demands mechanisms beyond pattern recognition, raising deeper questions about the nature of interaction and highlighting the importance of creativity in advancing conversational AI.
However, our analysis relies on the simplified single-layer self-attention setting from Li et al., 2024b . This does not directly extend to complex modern LLMs with deeper architectures and varied training goals. Future work will explore how these findings translate to more intricate models, examining their ability to change topics. Moreover, the recurring theme of topic continuity characterized in this work opens new avenues for applications, such as automatic change topic detection in conversational AI based on discrepancies between model predictions and speaker outputs.
@@ A.4 Proof of Theorem 5
Appendix B Detailed Explanation of Assumption 4
As illustrated in Figure 7, the dataset for $\mathbb{T}_{a}$ and the dataset for $\mathbb{T}_{b}$ demonstrate interchangeability with $\mathcal{G}_{a}^{(4)}$ and $\mathcal{G}_{b}^{(4)}$, respectively.
Figure 7: Illustration of Assumption 4. Here are two datasets related to the TPGs, $\mathcal{G}_{a}^{(4)}$ and $\mathcal{G}_{b}^{(4)}$, from Figure 3. From the directed arrows in $\mathcal{G}_{a}^{(4)}$, we can generate a dataset with the last token $\bf{e}_{4}$ for $\mathbb{T}_{a}$, which can reconstruct back the $\mathcal{G}_{a}^{(4)}$. A similar process applies for the $\mathcal{G}_{b}^{(4)}$.
Appendix C Detailed Simulation Studies
C.1 Simulation Process
@@
Next token prediction. To differentiate the input sequence length of the testing data from that of the training data, we introduce $T_{test}$. TPGs based on the training dataset DSETa are utilized to generate test datasets consisting of 100 sequences from $\mathbb{T}_{a}$ per epoch. Specifically, the last token ${\bf x}_{T_{test}}$ of the test input sequence is randomly selected from $K$ tokens (i.e. ${\bf x}_{T_{test}}={\bf e}_{k}$) and the remaining input tokens are randomly chosen based on the SCCs of $\mathcal{G}^{k}_{a}$, where tokens with higher priority are assigned greater weights. For instance, in $\mathcal{G}^{4}_{a}$, tokens ${\bf e}_{1},{\bf e}_{2},{\bf e}_{4}$ are captured with the priority order ${\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$. The weights assigned to input tokens ${\bf e}_{1},{\bf e}_{2}$, and ${\bf e}_{4}$ are $0.4,0.4$, and $0.2$, respectively. It reflects that ${\bf e}_{1}$ and ${\bf e}_{2}$ are in the same higher-priority SCC, thus having greater weights compared to ${\bf e}_{4}$. Intuitively, tokens within the same SCC are more likely to co-occur than those from different SCCs. This approach enables the generated test input sequences to mimic real word relationships and reflect their contextual groupings. Following the generation of the test dataset from $\mathbb{T}_{a}$, the next tokens $\hat{y}_{\bf w_{a}}$ and $\hat{y}_{\bf w_{ab}}$ are predicted by Equation 1, with ${\bf W}_{a}$ and ${\bf W}_{ab}$ obtained from the last iteration. To reduce the potential numerical issues in the outputs, $\mathbb{S}({\bf X}{\bf W}\bar{{\bf x}})$ is rounded to three decimals, ensuring that tokens within the same SCC yield consistent softmax outputs. Our code for all simulations is available in ( ).
C.2 Experiments Results
@@ C.2.1 Simulation in Section 4
C.2.2 Simulation in Section 5
@@
In Figure 6(a), we predict next tokens for 5000 test sequences from $\mathbb{T}_{a}$ with $T_{test}=\{4,8,16,24,32,64,128,256,512\}$, while fixing $L=4$, $d=16$, $T_{train}=4$, and $K=10$. The proportion of each scenario with varying $T$ is illustrated in Table 1. For Figure 6(b), we predict next tokens for 5000 test sequences (the sequence length is $T_{test}=20$) using models trained with $L=\{4,6,8,10,12,14,16,18\}$, $d=16$, $K=10$, and $T_{train}=4$. The proportion of each scenario with varying $L$ is illustrated in Table 2.
C.2.3 Additional experiments
@@
C.3 Numerical Analysis of each Scenario in Figure 4